# From Social Division to Cohesion with AI Message Suggestions in Online Chat Groups

### v1

@@ 2.3 From Conversational Receptiveness to Social Cohesion
Figure 7: Structural equation model of session-level effects of AI assistance on stance assortativity. Message volume, sentiment, and their interaction predict stance assortativity in the subsequent round, controlling for prior assortativity. The effects of individual and relational assistance are assessed relative to the no-assistance control condition. Solid arrows indicate standardized path coefficients, and dashed lines indicate covariances among mediators. Black lines indicate paths siginificant at $p<0.001$, and gray lines indicate paths that do not meet this threshold ($N=60$). Covariance estimates and indirect effects are provided in Extended Data Table 7.
3 Discussion
Our experiment demonstrates how LLM-powered message suggestions produce divergent effects on group composition dynamics in online discussions. Individual assistance, which tailored suggestions to users’ own stances, increased communication volume and group size but widened stance gaps across groups. Relational assistance, which incorporated awareness of other group members’ perspectives, fostered more emotionally positive and receptive exchanges, encouraging participants to form groups with greater opinion diversity. These contrasting pathways illustrate how subtle design choices in AI assistance can either entrench division or actively cultivate cohesion at scale.
Addressing social division and polarization is widely recognized as a pressing societal challenge [16, 1, 3, 29]. Prior work has tended to treat individual-level and system-level interventions separately [1, 41] with a limited understanding of how effects at one level propagate to the other. At the individual level, studies frequently focus on dyadic communication and examine how people might be guided toward consensus on controversial topics [42, 43, 27]. At the system level, interventions aim to increase opportunities for diverse populations to interact, such as by designing platforms that encourage cross-cutting exposure [32, 44]. However, these approaches remain disconnected: Individual-level studies lack mechanisms to evolve or even sustain opportunities for ongoing interaction [3], while system-level interventions often fall short because mere exposure is insufficient for social cohesion and can even backfire [32].
Our work complements these studies by explicitly addressing the dynamics between individual behavior and collective structure [6]. Rather than aiming for immediate consensus at the interpersonal level [29], we examine how the form of conversation can shape group communication norms and foster connections across differences [45]. Cohesion at the system level emerged through spontaneous interaction among individuals exercising free choice [46, 47]. Prior empirical work shows that certain linguistic patterns facilitate the formation of echo chambers online [48, 49]. Our findings suggest that targeted linguistic interventions can reverse this tendency, counteracting polarization dynamics and fostering more heterogeneous and cohesive group structures.
This study also highlights the risks and potential of AI assistance in shaping patterns of social division and cohesion. As AI technologies continue to advance, they have become embedded in everyday communication tools and are often personalized to individual users [21, 22]. While such personalization can improve usability, sycophantic AI assistance that echoes users’ views [26] may also amplify perceptual asymmetries, where “in-group virtues become out-group vices” [50].
This conventional personalization paradigm reflects methodological individualism in social science, which treats individuals as isolated units of analysis [51]. Our findings suggest complementing this perspective with a relational lens [52]. Individuals are not fixed entities but enact different roles and identities across social contexts [53, 54, 55]. If we view a “person” not only as an isolated individual but as someone embedded in interactional networks [56], AI personalization can be designed to help users express their opinions in ways that remain open and responsive to others—even those who disagree [57].
Human agency is another key consideration in the use of AI assistance. In our study, participants sent more messages when suggestions were available but produced fewer self-written ones, and they rarely edited the suggested content. These patterns suggest that participants delegated part of their agency to AI, primarily to reduce the effort of composing messages themselves [24]. As a result, relational assistance facilitated more receptive conversations and greater social cohesion, but these outcomes could occur without participants’ explicit intention. This dynamic resembles classic solutions to collective action problems, in which individual incentives help achieve public goods [58]. Our findings further suggest that humans may learn from AI by adopting more receptive language, as self-written messages became more emotionally positive when exposed to relational suggestions [27]. At the same time, however, the results raise concerns that human agency and interpersonal trust may erode as control over social interactions shifts toward automated systems [22, 59, 60, 61].
We acknowledge several limitations of this study that point to directions for future work. First, our experiment took place in a controlled online setting with structured discussions, whereas real-world conversations unfold over longer time horizons across multiple platforms [62]. Participants in our study rarely changed their opinion stance within a brief session, yet over extended interactions, established clusters may further polarize or, conversely, moderate through repeated exposure and trust-building [63, 64]. Second, future work can also advance the design space of conversational assistance itself. Although relational assistance effectively reduced stance-based assortativity, it was used less frequently than individual assistance, which led to higher conversational engagement. Exploring additional design dimensions—such as embodiment and controllability—may help balance social receptiveness with engagement [65, 66]. Finally, beyond technical refinement, evaluating these systems in politically salient contexts will be crucial for understanding how AI assistance operates amid the social and emotional complexities of real-world discourse [17, 15].
Taken together, this study demonstrates how subtle, AI-mediated changes in communication can scale from individual interactions to collective structures. By tracing the micro–macro loop, we show that message-level interventions can cascade into group-composition dynamics, either reinforcing divisions or fostering cohesion in opinion-diverse populations. As AI tools become increasingly ubiquitous, they offer an opportunity to redirect their influence toward the production of public goods. Reimagining personalization and autonomy as socially embedded processes may help resist the drift toward fragmentation and polarization that increasingly characterizes many societies today. Our findings suggest design levers for AI systems and communication policies that could facilitate more cohesive and resilient public discourse.
4 Methods
4.1 Ethics, Consent, and Preregistration
@@ 4.4 Message Suggestions
4.5 Analyses
4.5.1 Group Composition
@@ 4.5.2 Conversation Dynamics
4.5.3 Statistical Modeling from Conversation to Cohesion

### v2

@@ Results
Figure 7: Structural equation model of session-level effects of AI assistance on stance assortativity. Message volume, sentiment, and their interaction predict stance assortativity in the subsequent round, controlling for prior assortativity. The effects of individual and relational assistance are assessed relative to the no-assistance control condition. Solid arrows indicate standardized path coefficients, and dashed lines indicate covariances among mediators. Black lines indicate paths siginificant at $p<0.001$, and gray lines indicate paths that do not meet this threshold. Covariance estimates and indirect effects are provided in Extended Data Table 7.
Discussion
Subtle differences in AI-mediated expressive framing can propagate through conversational dynamics to reorganize the structure of online groups. Although assortativity is a session-level measure that usually shifts gradually, we observe clear divergence across conditions. Individual assistance, which tailored suggestions to users’ own stances, amplified communication volume yet increased separation between groups. Relational assistance, which incorporated awareness of others’ perspectives, fostered more receptive conversations and produced more heterogeneous, cross-cutting group configurations. These results indicate that AI-generated phrasing has become part of the interactional substrate through which social structure is produced.
Addressing the dynamics of social division and polarization remains a critical scientific and societal challenge [42, 15, 16, 30]. Research has largely bifurcated into two levels of analysis [15, 3]. Individual-level interventions aim to improve interpersonal receptiveness [43, 44, 27], whereas system-level interventions seek to modify exposure and contact patterns [33, 45]. Yet how effects at one level translate to the other remains challenging: interpersonal interventions seldom show how conversational shifts propagate into structural outcomes [16], and system-level interventions often falter because externally imposed exposure does not reliably sustain constructive interaction and can even exacerbate division [33].
Our study links these levels by showing how local expressive cues introduced during conversation can scale into collective structural patterns. Rather than targeting immediate persuasion [46, 47] or consensus [30], we focus on how the form of conversation—whether self-focused or socially aware—shapes the interactional norms that govern group formation and reorganization [48]. This complements prior work showing that specific linguistic styles can intensify ideological clustering [10, 11]. Our findings demonstrate the converse: when expressive framing highlights relational context, micro-level dynamics can instead dampen segregation pressures.
The results also illuminate properties of hybrid human–AI expressive systems [20, 22]. Personalization algorithms tend to mirror users’ existing stances [25, 26], potentially reinforcing perceptual asymmetries in which “in-group virtues become out-group vices” [49]. This paradigm reflects a form of methodological individualism [50], in which individuals are operationally treated in analytic isolation.
Our study points to a complementary relational perspective on AI framing [51, 52]. Individuals enact different roles, identities, and expectations across social contexts [53, 54, 55] and are situated within interactional networks [56]. Even modest changes in expressive framing—from “help me say what I think” to “help me say what I think in relation to them”—can shift how people engage and how groups subsequently configure themselves. Such AI framing can influence group dynamics even when only some participants make use of the suggestions, consistent with evidence that visible expressive cues can shape broader interactional norms [57, 58].
Human agency is another important consideration. In our study, participants sent more messages when suggestions were available but rarely edited them, suggesting partial delegation of expressive effort to the AI tool [24]. As a result, relational assistance yielded more cohesive structures, yet these outcomes did not require explicit intentions from participants. This dynamic resembles classic solutions to collective action problems, where selective incentives align individual behavior with group-level outcomes [59].
We also find evidence of expressive alignment through dual exposure effects: participants’ self-written messages became more emotionally positive and respectful when they were intermittently exposed both to relational AI suggestions and to sentiment-enhanced messages from group members [27]. This illustrates that hybrid expressive systems can shape not only message output but also users’ own linguistic patterns through social reinforcement. At the same time, increased reliance on automated support, even relational assistance, raises questions about how agency and interpersonal trust may shift as AI mediates a larger fraction of interactional work, and whether stylistic positivity or accommodation might mask substantive disagreement in more consequential settings [19, 60, 61, 62].
Several limitations suggest directions for future research. First, our experiment was conducted in a controlled online environment with short discussions, whereas real interactions unfold over longer time periods and across multiple platforms [63]. We interpret our results as revealing transferable interactional mechanisms rather than making direct predictions about full-scale online platforms. These mechanisms—shifts in conversational tone, asymmetries in message production, and the diffusion of hybrid human–AI expression—could plausibly operate in larger, persistent online communities, making this an important direction for future research. For example, in our study, participants rarely changed their own stance, but their opinions may polarize or moderate through longer, repeated exposure and relationship dynamics [64, 65]. Second, the design space of conversational assistance remains large. Although relational assistance reduced assortativity, it was used less frequently than individual assistance. Exploring additional forms of expressive support—such as timing, controllability, embodiment, or intent framing—may clarify how assistance can function more effectively as a socio-cognitive scaffold [47, 66, 67, 68]. Finally, deploying such systems in politically salient or emotionally charged contexts may reveal additional dynamics related to identity, conflict, and misinformation [69, 70].
This study provides an experimental test of the micro-macro consequences of AI-mediated expression. We show that small expressive shifts introduced by AI at the level of individual messages can propagate through conversational interaction to reshape collective outcomes. When AI assistance incorporates the social context surrounding users, it fosters more receptive communication and consequently reshapes social structure in more heterogeneous and cohesive ways within opinion-diverse populations. As AI systems become increasingly integrated into everyday communication, understanding how hybrid expressive processes diffuse through interaction offers a foundation for explaining—and anticipating—how local expressive cues reshape the macro-level organization of social systems.
Methods
Ethics, Consent, and Preregistration
@@ Methods
Analyses
Group Composition
@@
Statistical Modeling from Conversation to Cohesion